You have been hired as employee number 20 of a start-up company and have been offered...

  1. You have been hired as employee number 20 of a start-up company and have been offered options as a portion of your compensation. The options vest over 4 years, with 25% vesting at the end of the first year of your employment and the remaining 75% vesting monthly over the next three years. If you quit to join a new company halfway through your third year of employment (so you were employed for 2 ½ years), then what percentage of your options have vested? (to a tenth of a percent).

__________

  1. You have been granted options for 25,000 shares with an exercise price of $10 per share. If in the future you are fully vested, and if you cashed out your options for what they are worth on a net basis, what total net value did you receive if the stock price was $30 per share?
Answer #1

answer

1) percentage of option at the end of 2.5 years.

a) percentage of option vested at the end of 1 year= 25%

b) percentage of option vested for next 18 month= (75%/30 months)*18= 37.50%

so percentage of option vested at the end of 2.5 years= a+b

= 25%+37.50%= 62.50%

Explanation for b) Since question is saying 75% of option will vest in next 3 year every month. So determined percentage of vest every month.

answer 2) On net basis employee will receive $ 20*25000= $ 500000